# Robust Photo-Realistic Hand Gesture Generation: from Single View to Multiple View

### v1

@@ 4.4 Qualitative Results
Figure 8 illustrates the results of the qualitative analysis. It is shown that the hand regions generated by our proposed MUFEN are more realistic, and the gestures are more accurate.
5 Conclusion
@@
This paper proposed a diffusion model-based framework to enhance the realism accuracy of generated hand gestures in human-centric generation tasks through multi-view priors. Specifically, we employed coordinate transformations on 3D hand meshes combined with viewpoint camera position adjustments to derive six orthogonal mesh projections (front, back, left, right, top, bottom). Through quantitative evaluation of the projected area, we identified the optimal pair of viewpoints containing maximum geometric information for prior integration. We then designed a dedicated Multi-Modal UNet-based Feature Encoder (MUFEN) to fuse the complete hand modeling features from multi-view mesh and multi-modal features contained in other modalities and enhanced the accuracy of feature fusion with localization feature from the bounding box. Experiments demonstrated that the proposed framework significantly improved gesture generation quality, achieving SOTA performance across both quantitative metrics and qualitative assessments.

### v2

@@ 4.4. Qualitative Results
Figure 1 and Figure 7 illustrate the results of the qualitative analysis. It is shown that the hand regions generated by our proposed MUFEN are more realistic, and the gestures are more accurate.
4.5. Ablation Studies
Table 3 shows ablation results on different modalities and number of views. Clearly, MUFEN with original settings consistently outperforms all ablated variants on hand-region metrics, validating our design choices. Regarding the choice of number of views, using more predefined views (4 or 6) degrades performance, due to information redundancy from views overlap. While MUFEN selects the co-axial view pair with the largest combined projected area (2 views) as input, which minimizes redundancy and enhances gesture representation, achieving better performance and validating our complementary view selection strategy. Regarding the combination of multimodal, removing the depth map causes the most degradation (FID +8.60, FID-Hand +11.26), indicating its importance as a structural prior. Excluding the mesh also significantly affects performance (FID +4.58, FID-Hand +4.52), confirming its role in gesture geometry modeling. Gesture labels yield moderate gains (FID +1.17), while bounding box fusion helps localize hand regions, as reflected in increased FID-Hand (+0.88) when removed. These results affirm the importance of each modality and corresponding module in our framework and support the effectiveness of MUFEN’s multimodal design.
5. Conclusions
@@
This paper proposed a diffusion model-based framework to enhance the realism and accuracy of generated hand gestures in human-centric generation tasks through multi-view priors. Specifically, we employed coordinate transformations on 3D MANO mesh rendering combined with viewpoint camera position adjustments to derive six orthogonal mesh projections (front, rear, left, right, top, bottom). We identified the optimal pair of viewpoints that contain the maximum geometric information for prior integration via projected areas of the rendered meshes. We then designed a dedicated Multi-Modal UNet-based Feature Encoder (MUFEN) to fuse the comprehensive hand modeling features from multi-view meshes and multi-modal features contained in other modalities and enhanced the accuracy of feature fusion with localization feature from the bounding box. Experiments demonstrated that the proposed framework significantly improved gesture generation quality, achieving SOTA performance across both quantitative metrics and qualitative assessments.
6. Acknowledgments
This work was supported by the China Scholarship Council. It was also supported by Engineering and Physical Sciences Research Council [grant number EP/Y009800/1], through Keystone project funding from Responsible Ai UK (KP0016). This research utilized Queen Mary’s Andrena HPC facility, supported by QMUL Research-IT.